# Extremal correlation coefficient for functional data

### v2

@@ Appendix A Preliminary results
$E\left[f({\bf Z}_{n})\right]\leq\frac{\sqrt{n}}{k}\left\{\frac{k}{n}\frac{n}{k}E\left[\left(\frac{R_{1}}{b(n/k)}\right)^{2}I_{R_{1}\geq b(n/k)}\right]\right\}^{1/2}\leq\frac{c_{3}}{\sqrt{k}},$
@@
Proof of Lemma 5.1:  We begin by noting that since $Z_{1}$ and $Z_{2}$ are independent, there exists $\nu$ in $M_{+}({\mathbb{R}}_{+}^{2})$ such that
$n{\rm pr}\left(\frac{(|Z_{1}|,|Z_{2}|)}{b(n)}\in\cdot\right)\stackrel{{\scriptstyle\mbox{$\scriptstyle v$}}}{{\rightarrow}}\nu,$ (28)
and for ${\bf x}=[x_{1},x_{2}]^{\top}$
$\nu([0,{\bf x}]^{c})=c\{(x_{1})^{-\alpha}+(x_{2})^{-\alpha}\}.$
With the choice of $b(n)$ defined by
$n^{-1}$ $={\rm pr}(\|Z_{1}\phi_{1}\|\vee\|\rho Z_{1}\phi_{1}+\sqrt{1-\rho^{2}}Z_{2}\phi_{2}\|>b(n))$ (29)
$={\rm pr}(|Z_{1}|\vee(\rho^{2}Z_{1}^{2}+(1-\rho^{2})Z_{2}^{2})^{1/2}>b(n)),$
we set $c=1/(1+(1-\rho^{2})^{\alpha/2})$ to ensure that $\nu$ is a probability measure on $\{(z_{1},z_{2}):|z_{1}|\vee(\rho^{2}z_{1}^{2}+(1-\rho^{2})z_{2}^{2})^{1/2}>1\}$.
@@
$=\lim_{n\to\infty}E\left[\left\langle\frac{Z_{1}\phi_{1}}{b(n)},\frac{\rho Z_{1}\phi_{1}+\sqrt{1-\rho^{2}}Z_{2}\phi_{2}}{b(n)}\right\rangle\Bigg{|}\|Z_{1}\phi_{1}\|\vee\|\rho Z_{1}\phi_{1}+\sqrt{1-\rho^{2}}Z_{2}\phi_{2}\|>b(n)\right]$
$=\lim_{n\to\infty}\frac{1}{{\rm pr}(\|Z_{1}\phi_{1}\|\vee\|\rho Z_{1}\phi_{1}+\sqrt{1-\rho^{2}}Z_{2}\phi_{2}\|>b(n))}\times$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ E\left[\left\langle\frac{Z_{1}\phi_{1}}{b(n)},\frac{\rho Z_{1}\phi_{1}+\sqrt{1-\rho^{2}}Z_{2}\phi_{2}}{b(n)}\right\rangle I_{\|Z_{1}\phi_{1}\|\vee\|\rho Z_{1}\phi_{1}+\sqrt{1-\rho^{2}}Z_{2}\phi_{2}\|>b(n)}\right]$
$=\lim_{n\to\infty}\frac{1}{{\rm pr}(\|Z_{1}\phi_{1}\|\vee\|\rho Z_{1}\phi_{1}+\sqrt{1-\rho^{2}}Z_{2}\phi_{2}\|>b(n))}E\left[\rho\frac{Z_{1}^{2}}{b(n)^{2}}I_{|Z_{1}|\vee(\rho^{2}Z_{1}^{2}+(1-\rho^{2})Z_{2}^{2})^{1/2}>b(n)}\right]$
@@
$=\lim_{n\to\infty}E\left[\left\langle\frac{Z_{1}\phi_{1}}{b(n)},\frac{Z_{1}\phi_{1}}{b(n)}\right\rangle\Bigg{|}\|Z_{1}\phi_{1}\|\vee\|\rho Z_{1}\phi_{1}+\sqrt{1-\rho^{2}}Z_{2}\phi_{2}\|>b(n)\right]$
$=\lim_{n\to\infty}\frac{1}{{\rm pr}(\|Z_{1}\phi_{1}\|\vee\|\rho Z_{1}\phi_{1}+\sqrt{1-\rho^{2}}Z_{2}\phi_{2}\|>b(n))}E\left[\frac{Z_{1}^{2}}{b(n)^{2}}I_{|Z_{1}|\vee(\rho^{2}Z_{1}^{2}+(1-\rho^{2})Z_{2}^{2})^{1/2}>b(n)}\right]$
@@
$=\lim_{n\to\infty}E\left[\frac{\|\rho Z_{1}\phi_{1}+\sqrt{1-\rho^{2}}Z_{2}\phi_{2}\|^{2}}{b(n)^{2}}\Bigg{|}\|Z_{1}\phi_{1}\|\vee\|\rho Z_{1}\phi_{1}+\sqrt{1-\rho^{2}}Z_{2}\phi_{2}\|>b(n)\right]$
$=\lim_{n\to\infty}\frac{1}{{\rm pr}(\|Z_{1}\phi_{1}\|\vee\|\rho Z_{1}\phi_{1}+\sqrt{1-\rho^{2}}Z_{2}\phi_{2}\|>b(n))}\times$
@@
$=\int_{1}^{\infty}\rho^{2}z_{1}^{2}\ c\nu_{\alpha}(dz_{1})+\int_{1/(1-\rho^{2})^{1/2}}^{\infty}(1-\rho^{2})z_{2}^{2}\ c\nu_{\alpha}(dz_{1})$
=$=\left\{\rho^{2}+(1-\rho^{2})^{\alpha/2}\right\}\frac{c\alpha}{\alpha-2}.$
@@ D.1 Simulation results on the consistency of $\hat{\rho}_{n,k}$
This section reports the magnitude of empirical biases (the absolute difference between the average and the theoretical value), along with standard errors computed as the sample standard deviations. Using the optimal $k$s selected by the method from Danielsson et al. (2016), the results are shown in Tables 5 and 6 for $\alpha=\{4,5\}$. The results with the optimal $k$s selected by the method from Clauset et al. (2009) are provided in Tables 78, and 9, for $\alpha=\{3,4,5\}$.
@@ D.2 Relation to the measures $\chi$ and $\bar{\chi}$
If $U$ and $V$ are asymptotically independent, then $\bar{\chi}\in[-1,1)$, and if they are asymptotically dependent, then $\bar{\chi}=1$. These two measures are thus complementary and it is useful to apply them together. We will demonstrate that our ECC is complementary to them because it provides additional information on the shapes of the extremal curves.
@@ D.3 Effect of phase variation on $\rho_{XY}$
Phase variation occurs when some properties of curves shift over time, as seen with growth spurts occurring at different times for different individuals. Any variation in phase can affect the extremal correlation coefficient. In general, if extremal curves in the two samples are out of phase, the sample ECC will become closer to zero. For example, if heat waves tend arrive at different times at different locations, this will result in the ECC closer to zero than if the heat wave arrival times matched. To illustrate this, we consider the data generating process described in (17), but with $\phi_{k}$ in $Y(t)$, replaced by

### v3

@@ Appendix A Preliminary results
$E\left[f({\bf Z}_{n})\right]\leq\frac{n^{1/2}}{k}\left\{\frac{k}{n}\frac{n}{k}E\left[\left(\frac{R_{1}}{b(n/k)}\right)^{2}I_{R_{1}\geq b(n/k)}\right]\right\}^{1/2}\leq\frac{c_{3}}{k^{1/2}},$
@@
Proof of Lemma 5.1: We begin by noting that there exists an increasing sequence $b(n)\to\infty$ and $\nu$ in $M_{+}({\mathbb{R}}_{+}^{2})$ such that
$n{\rm pr}\left(\frac{(|Z_{1}|,|Z_{2}|)}{b(n)}\in\cdot\right)\stackrel{{\scriptstyle\mbox{$\scriptstyle v$}}}{{\rightarrow}}\nu.$ (28)
Here, $\nu$ is defined up to a multiplicative constant. Since $Z_{1}$ and $Z_{2}$ are independent, $\nu$ has its mass only on the axes.
The specific form of $\nu$ can be given by, for ${\bf x}=[z_{1},z_{2}]^{\top}$,
$\nu([0,{\bf x}]^{c})=c(z_{1})^{-\alpha}+c(z_{2})^{-\alpha},$ (29)
where $c=1/(1+(1-\rho^{2})^{\alpha/2})$. To see this, observe that
$n{\rm pr}(\|Z_{1}\phi_{1}\|\vee\|\rho Z_{1}\phi_{1}+(1-\rho^{2})^{1/2}Z_{2}\phi_{2}\|>b(n))=n{\rm pr}(|Z_{1}|\vee(\rho^{2}Z_{1}^{2}+(1-\rho^{2})Z_{2}^{2})^{1/2}>b(n)).$
By (5), it follows that $\nu(\{(z_{1},z_{2}):|z_{1}|\vee(\rho^{2}z_{1}^{2}+(1-\rho^{2})z_{2}^{2})^{1/2}>1\})$ must be 1. Note that $\{(z_{1},z_{2}):|z_{1}|\vee(\rho^{2}z_{1}^{2}+(1-\rho^{2})z_{2}^{2})^{1/2}>1\}=\{(z_{1},z_{2}):z_{1}>1\ {\rm or}\ z_{2}>1/(1-\rho^{2})^{1/2}\}$. Using this, we have that
$\nu(\{(z_{1},z_{2}):|z_{1}|\vee(\rho^{2}z_{1}^{2}+(1-\rho^{2})z_{2}^{2})^{1/2}>1\})=$ $c+c(1-\rho^{2})^{\alpha/2}=1.$
@@
$=\lim_{n\to\infty}E\left[\left\langle\frac{Z_{1}\phi_{1}}{b(n)},\frac{\rho Z_{1}\phi_{1}+(1-\rho^{2})^{1/2}Z_{2}\phi_{2}}{b(n)}\right\rangle\Bigg|\|Z_{1}\phi_{1}\|\vee\|\rho Z_{1}\phi_{1}+(1-\rho^{2})^{1/2}Z_{2}\phi_{2}\|>b(n)\right]$
$=\lim_{n\to\infty}\frac{1}{{\rm pr}(\|Z_{1}\phi_{1}\|\vee\|\rho Z_{1}\phi_{1}+(1-\rho^{2})^{1/2}Z_{2}\phi_{2}\|>b(n))}\times$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ E\left[\left\langle\frac{Z_{1}\phi_{1}}{b(n)},\frac{\rho Z_{1}\phi_{1}+(1-\rho^{2})^{1/2}Z_{2}\phi_{2}}{b(n)}\right\rangle I_{\|Z_{1}\phi_{1}\|\vee\|\rho Z_{1}\phi_{1}+(1-\rho^{2})^{1/2}Z_{2}\phi_{2}\|>b(n)}\right]$
$=\lim_{n\to\infty}\frac{1}{{\rm pr}(\|Z_{1}\phi_{1}\|\vee\|\rho Z_{1}\phi_{1}+(1-\rho^{2})^{1/2}Z_{2}\phi_{2}\|>b(n))}E\left[\rho\frac{Z_{1}^{2}}{b(n)^{2}}I_{|Z_{1}|\vee(\rho^{2}Z_{1}^{2}+(1-\rho^{2})Z_{2}^{2})^{1/2}>b(n)}\right].$
@@
$=\lim_{n\to\infty}E\left[\left\langle\frac{Z_{1}\phi_{1}}{b(n)},\frac{Z_{1}\phi_{1}}{b(n)}\right\rangle\Bigg|\|Z_{1}\phi_{1}\|\vee\|\rho Z_{1}\phi_{1}+(1-\rho^{2})^{1/2}Z_{2}\phi_{2}\|>b(n)\right]$
$=\lim_{n\to\infty}\frac{1}{{\rm pr}(\|Z_{1}\phi_{1}\|\vee\|\rho Z_{1}\phi_{1}+(1-\rho^{2})^{1/2}Z_{2}\phi_{2}\|>b(n))}E\left[\frac{Z_{1}^{2}}{b(n)^{2}}I_{|Z_{1}|\vee(\rho^{2}Z_{1}^{2}+(1-\rho^{2})Z_{2}^{2})^{1/2}>b(n)}\right]$
@@
$=\lim_{n\to\infty}E\left[\frac{\|\rho Z_{1}\phi_{1}+(1-\rho^{2})^{1/2}Z_{2}\phi_{2}\|^{2}}{b(n)^{2}}\Bigg|\|Z_{1}\phi_{1}\|\vee\|\rho Z_{1}\phi_{1}+(1-\rho^{2})^{1/2}Z_{2}\phi_{2}\|>b(n)\right]$
$=\lim_{n\to\infty}\frac{1}{{\rm pr}(\|Z_{1}\phi_{1}\|\vee\|\rho Z_{1}\phi_{1}+(1-\rho^{2})^{1/2}Z_{2}\phi_{2}\|>b(n))}\times$
@@
$=\int_{1}^{\infty}\rho^{2}z_{1}^{2}\ c\nu_{\alpha}(dz_{1})+\int_{1/(1-\rho^{2})^{1/2}}^{\infty}(1-\rho^{2})z_{2}^{2}\ c\nu_{\alpha}(dz_{2})$
$=\rho^{2}\frac{c\alpha}{\alpha-2}+(1-\rho^{2})^{\alpha/2}\frac{c\alpha}{\alpha-2}=\{\rho^{2}+(1-\rho^{2})^{\alpha/2}\}\frac{c\alpha}{\alpha-2}.$
@@
An alternative approach to proving Lemma 5.1 is to identify the joint angular measure $\Gamma$ of $X,Y]^{\top}$ and then apply it in the decomposition given by (11). Since only one of $Z_{1}$, $Z_{2}$ can be extreme, $\Gamma$ concentrates on $(\phi_{1},\rho\phi_{1})$ with mass $\nu(\{(z_{1},z_{2}):|z_{1}|\vee|\rho z_{1}|>1\})=c$ or $(\mbox{${0}$},\phi_{2})$ with mass $\nu(\{(z_{1},z_{2}):0\vee|(1-\rho^{2})^{1/2}z_{2}|>1\})=c(1-\rho^{2})^{\alpha/2}$. Therefore, we have that
$\Gamma=c\epsilon_{(\phi_{1},\rho\phi_{1})}+c(1-\rho^{2})^{\alpha/2}\epsilon_{(\mbox{${0}$},\phi_{2})}.$
Computing $\Gamma$ in (11), we obtain (30). Similarly, (31) and (32) follow.
@@ D.1 Simulation results on the consistency of $\hat{\rho}_{n,k}$
This section reports the magnitude of empirical biases, measured by the absolute difference between the average and the theoretical value, along with standard errors computed as the sample standard deviations. Using the optimal $k$s selected by the method from Danielsson et al. (2016), the results are shown in Tables 5 and 6 for $\alpha=\{4,5\}$. The results with the optimal $k$s selected by the method from Clauset et al. (2009) are provided in Tables 78, and 9, for $\alpha=\{3,4,5\}$.
@@ D.2 Relation to the measures $\chi$ and $\bar{\chi}$
If $U$ and $V$ are asymptotically independent, then $\bar{\chi}\in[-1,1)$, and if they are asymptotically dependent, then $\bar{\chi}=1$. These two measures are thus complementary, and it is useful to apply them together. We will demonstrate that our extremal correlation coefficient is complementary to them because it provides additional information on the shapes of the extremal curves.
@@ D.3 Effect of phase variation on $\rho_{XY}$
Phase variation occurs when some properties of curves shift over time, as seen with growth spurts occurring at different times for different individuals. Any variation in phase can affect the extremal correlation coefficient. In general, if extremal curves in the two samples are out of phase, the sample extremal correlation coefficient will become closer to zero. For example, if heat waves tend to arrive at different times at different locations, this will result in the extremal correlation coefficient closer to zero than if the heat wave arrival times matched. To illustrate this, we consider the data generating process described in (17), but with $\phi_{k}$ in $Y(t)$, replaced by